# Tutorial on amortized optimization

### v4

@@ 2.3.4 Settings with additional learnable contexts $\varphi$
where $y^{\star}(x,\varphi)$ can be replaced with an approximation by a learned amortization model $\hat{y}_{\theta}\approx y^{\star}$. The parameters $\varphi$ in eq. 2.28 can often by end-to-end learned through the solution of eq. 1.1 to update the influence that $\varphi$ has on the solutions. The next section turns to methods that show how to differentiate through the value $f(y^{\star}(x,\varphi);x,\varphi)$ and solution $y^{\star}(x,\varphi)$ to enable gradient-based learning of $\varphi$ in eq. 2.28.
@@
Methods can end-to-end learn through the optimal objective value $f(y^{\star}(x,\varphi);x,\varphi)$ to update parameters $\varphi$ that show up in the context — i.e. by taking $\ell=f$ in eq. 2.28. For example, variational autoencoders differentiate through the objective value, i.e. the best approximation to the ELBO, to optimize the data log-likelihood of a parameterized decoder $\log p(x\mid z;\varphi)$. The theory around this is rooted in the optimization community’s studies of envelope theorems, which describe settings where the minimum value can be differentiated by just differentiating the objective. Danskin’s envelope theorem [Danskin, 1966] in convex settings is one of the earliest and has been extended into more general settings, e.g., in [Bertsekas, 1971, Prop. A.22] and Carter [2001], Milgrom and Segal [2002], Bonnans and Shapiro [2013]. In the unconstrained and non-convex eq. 1.1, the envelope theorem gives
@@
Computing the adjoint derivative $\mathrm{D}_{\varphi}y^{\star}(x,\varphi)$ is more involved than the value derivative using the envelope theorem in eq. 2.29 as more components of $y^{\star}(x)$ can change as $x$ moves infinitesimally. An explicit closed-form solution to $y^{\star}(x)$ is not available in most cases, which means that standard methods for explicitly computing the derivative through this computation may not work well or may break down. For example, an optimizer to compute $y^{\star}(x)$ may be explicitly unrolled through, but this may be unstable and extremely memory- and compute-intensive to track all of the iterations. The adjoint derivative is typically computed with implicit differentiation by seeing $y^{\star}(x)$ as an implicit function of $x$. This uses the implicit function theorem, which is originally from Dini [1878], and is presented in Dontchev and Rockafellar [2009, Theorem 1B.1] as:
@@
The adjoint derivative $D_{\varphi}y^{\star}(\varphi)$ can be computed by seeing $y^{\star}$ as the root of an implicit function $g(y;x,\varphi)$, which needs to be selected to make the solution equivalent to the solution of eq. 1.1. Typically $g(y;x,\varphi)$ is an optimality system of the optimization problem, e.g. the KKT system for constrained convex optimization problems. For the unconstrained problem here, the first-order optimality of the objective $g(y;x,\varphi)\vcentcolon=\nabla_{y}f(y;x,\varphi)$ can be used with theorem 1 to compute $\mathrm{D}_{\varphi}y^{\star}(x,\varphi)$.
@@
Table 3.1: Applications of amortized optimization covered in chapter 3
@@
3.1 VAE $-\operatorname{ELBO}$ variational posterior data full ${\mathcal{L}}_{\rm obj}$
@@ 3.5 Meta OT dual OT cost optimal couplings input measures full ${\mathcal{L}}_{\rm obj}$
3.6 BC/IL $-Q$-value controls state space full ${\mathcal{L}}_{\rm reg}$
@@
SVG $\mathrm{D}_{\mathcal{Q}}$ or $-{\mathcal{E}}_{Q}$ control dists | full ${\mathcal{L}}_{\rm obj}$
@@
This section takes a review and tour of many key applications of amortized optimization to show some unifying ideas that can be shared between all of these topics. Table 3.1 summarizes the methods. The subsections in here are meant to be standalone and can be randomly accessed and read in any order. I scope closely to providing the relevant context for just the amortized optimization components and under-emphasize the remaining context of each research area.
@@ 3.1.1 The variational autoencoder (VAE) by Kingma and Welling [2014]
[…]
@@
Gaussians are a common choice of the variational distribution $q(z;\lambda)$ is in Kingma and Welling [2014], but may cause a loose inequality in eq. 3.2. Rezende and Mohamed [2015], Cremer et al. [2018] explore more expressive distributions to help make $\operatorname{ELBO}(\lambda^{\star};x,\varphi)$ equal to $\log p(x)$.
Amortized VI (AVI) methods predict the solution to eq. 3.3 while stochastic variational methods [Hoffman et al., 2013] explicitly solve it. AVI methods learn a model $\hat{\lambda}_{\theta}:{\mathcal{X}}\rightarrow\Lambda$ with parameters $\theta$, which is usually a feedforward neural network, to predict the maximum value of the $\operatorname{ELBO}$ by optimizing the objective-based loss
@@
Summary. This standard AVI formulation is therefore an amortized optimization method ${\mathcal{A}}_{\rm VAE}\vcentcolon=(-\operatorname{ELBO},\Lambda,{\mathcal{X}},p(x),\hat{\lambda}_{\theta},{\mathcal{L}}_{\rm obj})$ over the (negated) $\operatorname{ELBO}$ where the domain of the optimization problem is the variational parameter space $\Lambda$, the context space ${\mathcal{X}}$ is the sample space for the generative model, the samples are given from $p(x)$, $\hat{\lambda}_{\theta}:{\mathcal{X}}\rightarrow\Lambda$ is the fully-amortized model optimized with the gradient-based loss ${\mathcal{L}}_{\rm obj}$ over $-\operatorname{ELBO}$.
Extensions. Analyzing and extending the amortization components has been a key development in AVI methods. Cremer et al. [2018] investigate suboptimality in these models and categorize it as coming from an amortization gap where the amortized model for eq. 3.4 does not properly solve it, or the approximation gap where the variational posterior is incapable of approximating the true distribution. Semi-amortization plays a crucial role in addressing the amortization gap and is explored in the semi-amortized VAE (SAVAE) by Kim et al. [2018] and iterative VAE (IVAE) by Marino et al. [2018b]. AVI methods are also used in hierarchical [Sønderby et al., 2016] and sequential settings [Chung et al., 2015].
@@ 3.2 Sparse coding
where $E$ is the regularized reconstruction energy and $\alpha\in\mathbb{R}_{+}$ is a coefficient of the $\ell_{1}$ term. Eq. 3.6 is traditionally solved with the Iterative Shrinkage and Thresholding Algorithm (ISTA) such as in Daubechies et al. [2004]. Fast ISTA (FISTA) by Beck and Teboulle [2009] improves ISTA even more by adding a momentum term. The core update of ISTA methods is
@@
$W_{e}\vcentcolon=(1/L)W_{d}^{\top}$ is the filter matrix, $L$ is an upper bound on the largest eigenvalue of $W_{d}^{T}W_{d}$, $S\vcentcolon=I-W_{e}W_{d}$ is the mutual inhibition matrix, and $h_{\beta}(v)\vcentcolon=\operatorname{sign}(v)\max\left\{0,|v|-\beta\right\}$ is the shrinkage function with threshold $\beta$, usually set to $\alpha/L$. ISTA methods are remarkably fast ways of solving eq. 3.6 and the machine learning community has explored the use of learning to make ISTA methods even faster that can be seen as instances of amortized optimization.
@@ 3.2.2 Learned ISTA (LISTA) by Gregor and LeCun [2010]
LISTA further explores the idea of predicting solutions to sparse coding problems by proposing a semi-amortized model that integrates the iterative updates of ISTA into the model. LISTA leverages the soft-thresholding operator $h$ and considers a semi-amortized model over the domain ${\mathcal{Y}}$ that starts with $x^{0}_{\theta}\vcentcolon=0$ and iteratively updates $x^{t+1}_{\theta}\vcentcolon=h_{\beta}(Fx+Gx_{\theta}^{t})$. Running these updates for $K$ steps results in a final prediction $\hat{y}(x)\vcentcolon=x^{K}_{\theta}$ parameterized by $\theta=\{F,G,\beta\}$ that is also optimized with the regression-based loss to the ground-truth sparse codes as in eq. 3.9.
@@ 3.3 Multi-task learning and meta-learning
The motivation here is that there is likely shared structure and information present between the tasks that learning methods can leverage. The next section goes through methods that solve eq. 3.10 using objective-based amortized optimization methods.
@@ 3.3.2 Learning to optimize (LM) by Li and Malik [2017a]
As discussed in section 2.1.2, MAML can be seen as a semi-amortized optimization method. They also seek to predict the parameters $\hat{\varphi}_{\theta}({\mathcal{T}})\in\Phi$ of prediction model $\hat{y}_{\hat{\varphi}_{\theta}({\mathcal{T}})}(x)=\vcentcolon\hat{y}_{\theta}(x)$ in a multi-task setting with tasks ${\mathcal{T}}\sim p({\mathcal{T}})$. They propose to only learn to predict an initial iterate $\hat{\varphi}^{0}_{\theta}({\mathcal{T}})=\theta$ and then update the next iterates with gradient-based updates such as
@@ 3.3.4 Protein MSA modeling with the Neural Potts Model
Sercu et al. [2021] proposes a fully-amortized solution to fit a Potts model to a protein’s multiple sequence alignment (MSA). Each task consists of a finite MSA :=${\mathcal{M}}\vcentcolon=\{x_{i}\}$ and they use a fully-amortized model $\varphi_{\theta}({\mathcal{M}})\in\Phi$ to predict the optimal parameters of a Potts model $p({\mathcal{M}};\varphi)$ fit to the data. The model $\varphi_{\theta}$ is a large attention-based sequence model that takes the MSA as the input. Learning is done with the objective-based loss
@@ 3.3.5 Other relevant multi-task and meta-learning work
    Metz et al. [2019a] study semi-amortized optimization of convolutional architectures and identify and focus on key issues of 1) biased gradients from truncated BPTT and 2) exploding gradient norms from unrolling for many timesteps. They overcome both of these issues by optimizing the smoothed loss in eq. 2.14 with a variant of the gradient estimator proposed in Parmas et al. [2018] for reinforcement learning. This estimator re-weights reparameterization gradients and likelihood ratio gradients using inverse variance weighting [Fleiss, 1993]. Parmas and Sugiyama [2021] further unify the likelihood ratio and reparameterization gradients by connecting them with the divergence theorem which enables them to create a generalized estimator combining them.
@@ 3.4 Fixed-point computations and convex optimization
Continuous fixed-point problems as in definition 5 and illustrated in fig. 3.1 are ubiquitous in engineering and science and amortizing their solutions is an activate research area. Let ${\mathcal{R}}(y;x)\vcentcolon=g(y;x)-y$ be the fixed-point residual with squared norm ${\mathcal{N}}(y;x)\vcentcolon=\|{\mathcal{R}}(y;x)\|_{2}^{2}$. Fixed-point computations are connected to continuous unconstrained optimization as any fixed-point problem can be transformed into eq. 1.1 by optimizing the residual norms with:
@@
and conversely eq. 1.1 can be transformed into a fixed-point problem via first-order optimality to find the fixed-point of $\nabla f(y;x)-y=0$. Thus methods that to amortize the solutions to eq. 1.1 can help amortize the solutions to fixed-point computations definition 5.
@@
where $\beta\in[0,1]$ is a coefficient that controls whether the iterates or application of $g$ on the iterates should be used, and $\alpha\in\mathbb{R}^{M+1}$ where $11^{\top}\alpha=$ 1 are the coefficients used to combine the previous iterates. A basic AA method sets $\beta=1$ and solves
@@
using an acceleration model ${\rm acc_{\theta}}$ that is typically a recurrent model that predicts the next iterate given the past sequence of iterates. Venkataraman and Amos [2021, Prop. 1] discuss how this model captures standard AA as an instance by setting the models equal to the standard update that doesn’t use learning. They learn this model to amortize eq. 3.15 over a distribution of contexts $p(x)$ with an objective-based loss that solves
@@
[…]
where $x\in\mathbb{R}^{n}$ is the primal variable, $s\in\mathbb{R}^{m}$ is the primal slack variable, $y\in\mathbb{R}^{m}$ is the dual variable, and $r\in\mathbb{R}^{n}$ is the dual residual. The set $\mathcal{K}\in\mathbb{R}^{m}$ is a non-empty convex cone with dual cone $\mathcal{K}^{*}\in\mathbb{R}^{m}$. where $x\in\mathbb{R}^{n}$ is the primal variable, $s\in\mathbb{R}^{m}$ is the primal slack variable, $y\in\mathbb{R}^{m}$ is the dual variable, and $r\in\mathbb{R}^{n}$ is the dual residual. The set $\mathcal{K}\in\mathbb{R}^{m}$ is a non-empty convex cone. SCS uses the homogeneous self-dual embedding to view eq. 3.21 as a fixed-point computation over ${\mathcal{Z}}=\mathbb{R}^{n\times m\times 1}$ with a scalar-valued scaling factor as the last component.
NeuralSCS proposes a semi-amortized model to predict the solution to the fixed point of the self-dual embedding that solves eq. 3.21. Their semi-amortized model $\hat{z}_{\theta}(\phi)$ takes a context $\phi$ as the input and outputs a solution to the self-dual embedding by composing the SCS iterations $f$ with the learned fixed-point acceleration modules (${\rm init}_{\theta},{\rm acc}_{\theta}$).

### v5

@@ 2.3.4 Settings with additional learnable contexts $\varphi$
where $y^{\star}(x,\varphi)$ can be replaced with an approximation by a learned amortization model $\hat{y}_{\theta}\approx y^{\star}$. The parameters $\varphi$ in eq.˜2.28 can often by end-to-end learned through the solution of eq.˜1.1 to update the influence that $\varphi$ has on the solutions. The next section turns to methods that show how to differentiate through the value $f(y^{\star}(x,\varphi);x,\varphi)$ and solution $y^{\star}(x,\varphi)$ to enable gradient-based learning of $\varphi$ in eq.˜2.28.
@@
Methods can end-to-end learn through the optimal objective value $f(y^{\star}(x,\varphi);x,\varphi)$ to update parameters $\varphi$ that show up in the context — i.e. by taking $\ell=f$ in eq.˜2.28. For example, variational autoencoders differentiate through the objective value, i.e. the best approximation to the ELBO, to optimize the data log-likelihood of a parameterized decoder $\log p(x\mid z;\varphi)$. The theory around this is rooted in the optimization community’s studies of envelope theorems, which describe settings where the minimum value can be differentiated by just differentiating the objective. Danskin’s envelope theorem [Danskin, 1966] in convex settings is one of the earliest and has been extended into more general settings, e.g., in [Bertsekas, 1971, Prop. A.22] and Carter [2001], Milgrom and Segal [2002], Bonnans and Shapiro [2013]. In the unconstrained and non-convex eq.˜1.1, the envelope theorem gives
@@
Computing the adjoint derivative $\mathrm{D}_{\varphi}y^{\star}(x,\varphi)$ is more involved than the value derivative using the envelope theorem in eq.˜2.29 as more components of $y^{\star}(x)$ can change as $x$ moves infinitesimally. An explicit closed-form solution to $y^{\star}(x)$ is not available in most cases, which means that standard methods for explicitly computing the derivative through this computation may not work well or may break down. For example, an optimizer to compute $y^{\star}(x)$ may be explicitly unrolled through, but this may be unstable and extremely memory- and compute-intensive to track all of the iterations. The adjoint derivative is typically computed with implicit differentiation by seeing $y^{\star}(x)$ as an implicit function of $x$. This uses the implicit function theorem, which is originally from Dini [1878], and is presented in Dontchev and Rockafellar [2009, Theorem 1B.1] as:
@@
The adjoint derivative $D_{\varphi}y^{\star}(\varphi)$ can be computed by seeing $y^{\star}$ as the root of an implicit function $g(y;x,\varphi)$, which needs to be selected to make the solution equivalent to the solution of eq.˜1.1. Typically $g(y;x,\varphi)$ is an optimality system of the optimization problem, e.g. the KKT system for constrained convex optimization problems. For the unconstrained problem here, the first-order optimality of the objective $g(y;x,\varphi)\vcentcolon=\nabla_{y}f(y;x,\varphi)$ can be used with theorem˜1 to compute $\mathrm{D}_{\varphi}y^{\star}(x,\varphi)$.
@@
Table 3.1: Applications of amortized optimization covered in chapter˜3
@@
3.1 VAE $\operatorname{ELBO}$ variational posterior data full ${\mathcal{L}}_{\rm obj}$
@@ 3.5 Meta OT dual OT cost optimal couplings input measures full ${\mathcal{L}}_{\rm obj}$
3.6 BC/IL Q-Q-value controls state space full ${\mathcal{L}}_{\rm reg}$
@@
SVG $\mathrm{D}_{\mathcal{Q}}$ or ${\mathcal{E}}_{Q}$ control dists | full ${\mathcal{L}}_{\rm obj}$
@@
This section takes a review and tour of many key applications of amortized optimization to show some unifying ideas that can be shared between all of these topics. Table˜3.1 summarizes the methods. The subsections in here are meant to be standalone and can be randomly accessed and read in any order. I scope closely to providing the relevant context for just the amortized optimization components and under-emphasize the remaining context of each research area.
@@ 3.1.1 The variational autoencoder (VAE) by Kingma and Welling [2014]
A VAE models a density $p(x)$ over a high-dimensional space, for example images, text, or videos, given samples $x\sim p(x)$. They introduce a lower-dimensional latent space with a known distribution $p(z)$, such as an isotropic Gaussian, designed to capture hidden structure present in $p(x)$. VAEs parameterize a likelihood model $p(x;\varphi)$ with $\varphi$. Optimizing the log-likelihood $\log p(x;\varphi)=\log\int_{z}p(x\mid z;\varphi)p(z)dz$ with this latent structure is typically intractable because of the integral over the latent space. Variational methods overcome this by introducing a tractable lower-bound called the evidence lower bound ($\operatorname{ELBO}$) defined by
@@
Gaussians are a common choice of the variational distribution $q(z;\lambda)$ is in Kingma and Welling [2014], but may cause a loose inequality in eq.˜3.2. Rezende and Mohamed [2015], Cremer et al. [2018] explore more expressive distributions to help make $\operatorname{ELBO}(\lambda^{\star};x,\varphi)$ equal to $\log p(x)$.
Amortized VI (AVI) methods predict the solution to eq.˜3.3 while stochastic variational methods [Hoffman et al., 2013] explicitly solve it. AVI methods learn a model $\hat{\lambda}_{\theta}:{\mathcal{X}}\rightarrow\Lambda$ with parameters $\theta$, which is usually a feedforward neural network, to predict the maximum value of the $\operatorname{ELBO}$ by optimizing the objective-based loss
@@
Summary. This standard AVI formulation is therefore an amortized optimization method ${\mathcal{A}}_{\rm VAE}\vcentcolon=(-\operatorname{ELBO},\Lambda,{\mathcal{X}},p(x),\hat{\lambda}_{\theta},{\mathcal{L}}_{\rm obj})$ over the (negated) $\operatorname{ELBO}$ where the domain of the optimization problem is the variational parameter space $\Lambda$, the context space $\mathcal{X}$} is the sample space for the generative model, the samples are given from $p(x)$, $\hat{\lambda}_{\theta}:{\mathcal{X}}\rightarrow\Lambda$ is the fully-amortized model optimized with the gradient-based loss ${\mathcal{L}}_{\rm obj}$ over $\operatorname{ELBO}$.
Extensions. Analyzing and extending the amortization components has been a key development in AVI methods. Cremer et al. [2018] investigate suboptimality in these models and categorize it as coming from an amortization gap where the amortized model for eq.˜3.4 does not properly solve it, or the approximation gap where the variational posterior is incapable of approximating the true distribution. Semi-amortization plays a crucial role in addressing the amortization gap and is explored in the semi-amortized VAE (SAVAE) by Kim et al. [2018] and iterative VAE (IVAE) by Marino et al. [2018b]. AVI methods are also used in hierarchical [Sønderby et al., 2016] and sequential settings [Chung et al., 2015].
@@ 3.2 Sparse coding
where $E$ is the regularized reconstruction energy and $\alpha\in\mathbb{R}_{+}$ is a coefficient of the $\ell_{1}$ term. Eq.˜3.6 is traditionally solved with the Iterative Shrinkage and Thresholding Algorithm (ISTA) such as in Daubechies et al. [2004]. Fast ISTA (FISTA) by Beck and Teboulle [2009] improves ISTA even more by adding a momentum term. The core update of ISTA methods is
@@
$W_{e}\vcentcolon=(1/L)W_{d}^{\top}$ is the filter matrix, $L$ is an upper bound on the largest eigenvalue of $W_{d}^{T}W_{d}$, $S\vcentcolon=I-W_{e}W_{d}$ is the mutual inhibition matrix, and $h_{\beta}(v)\vcentcolon=\operatorname{sign}(v)\max\left\{0,|v|-\beta\right\}$ is the shrinkage function with threshold $\beta$, usually set to $\alpha/L$. ISTA methods are remarkably fast ways of solving eq.˜3.6 and the machine learning community has explored the use of learning to make ISTA methods even faster that can be seen as instances of amortized optimization.
@@ 3.2.2 Learned ISTA (LISTA) by Gregor and LeCun [2010]
LISTA further explores the idea of predicting solutions to sparse coding problems by proposing a semi-amortized model that integrates the iterative updates of ISTA into the model. LISTA leverages the soft-thresholding operator $h$ and considers a semi-amortized model over the domain $\mathcal{Y}$} that starts with $x^{0}_{\theta}\vcentcolon=0$ and iteratively updates $x^{t+1}_{\theta}\vcentcolon=h_{\beta}(Fx+Gx_{\theta}^{t})$. Running these updates for $K$ steps results in a final prediction $\hat{y}(x)\vcentcolon=x^{K}_{\theta}$ parameterized by $\theta=\{F,G,\beta\}$ that is also optimized with the regression-based loss to the ground-truth sparse codes as in eq.˜3.9.
@@ 3.3 Multi-task learning and meta-learning
The motivation here is that there is likely shared structure and information present between the tasks that learning methods can leverage. The next section goes through methods that solve eq.˜3.10 using objective-based amortized optimization methods.
@@ 3.3.2 Learning to optimize (LM) by Li and Malik [2017a]
As discussed in section˜2.1.2, MAML can be seen as a semi-amortized optimization method. They also seek to predict the parameters $\hat{\varphi}_{\theta}({\mathcal{T}})\in\Phi$ of prediction model $\hat{y}_{\hat{\varphi}_{\theta}({\mathcal{T}})}(x)=\vcentcolon\hat{y}_{\theta}(x)$ in a multi-task setting with tasks ${\mathcal{T}}\sim p({\mathcal{T}})$. They propose to only learn to predict an initial iterate $\hat{\varphi}^{0}_{\theta}({\mathcal{T}})=\theta$ and then update the next iterates with gradient-based updates such as
@@ 3.3.4 Protein MSA modeling with the Neural Potts Model
Sercu et al. [2021] proposes a fully-amortized solution to fit a Potts model to a protein’s multiple sequence alignment (MSA). Each task consists of a finite MSA ${\mathcal{M}}\vcentcolon=\{x_{i}\}$ and they use a fully-amortized model $\varphi_{\theta}({\mathcal{M}})\in\Phi$ to predict the optimal parameters of a Potts model $p({\mathcal{M}};\varphi)$ fit to the data. The model $\varphi_{\theta}$ is a large attention-based sequence model that takes the MSA as the input. Learning is done with the objective-based loss
@@ 3.3.5 Other relevant multi-task and meta-learning work
    Metz et al. [2019a] study semi-amortized optimization of convolutional architectures and identify and focus on key issues of 1) biased gradients from truncated BPTT and 2) exploding gradient norms from unrolling for many timesteps. They overcome both of these issues by optimizing the smoothed loss in eq.˜2.14 with a variant of the gradient estimator proposed in Parmas et al. [2018] for reinforcement learning. This estimator re-weights reparameterization gradients and likelihood ratio gradients using inverse variance weighting [Fleiss, 1993]. Parmas and Sugiyama [2021] further unify the likelihood ratio and reparameterization gradients by connecting them with the divergence theorem which enables them to create a generalized estimator combining them.
@@ 3.4 Fixed-point computations and convex optimization
Continuous fixed-point problems as in definition˜5 and illustrated in fig.˜3.1 are ubiquitous in engineering and science and amortizing their solutions is an activate research area. Let ${\mathcal{R}}(y;x)\vcentcolon=g(y;x)-y$ be the fixed-point residual with squared norm ${\mathcal{N}}(y;x)\vcentcolon=\|{\mathcal{R}}(y;x)\|_{2}^{2}$. Fixed-point computations are connected to continuous unconstrained optimization as any fixed-point problem can be transformed into eq.˜1.1 by optimizing the residual norms with:
@@
and conversely eq.˜1.1 can be transformed into a fixed-point problem via first-order optimality to find the fixed-point of $\nabla f(y;x)-y=0$. Thus methods that to amortize the solutions to eq.˜1.1 can help amortize the solutions to fixed-point computations definition˜5.
@@
where $\beta\in[0,1]$ is a coefficient that controls whether the iterates or application of $g$ on the iterates should be used, and $\alpha\in\mathbb{R}^{M+1}$ where $1^{\top}\alpha=1$ are the coefficients used to combine the previous iterates. A basic AA method sets $\beta=1$ and solves
@@
using an acceleration model $\rm acc_{\theta}$} that is typically a recurrent model that predicts the next iterate given the past sequence of iterates. Venkataraman and Amos [2021, Prop. 1] discuss how this model captures standard AA as an instance by setting the models equal to the standard update that doesn’t use learning. They learn this model to amortize eq.˜3.15 over a distribution of contexts $p(x)$ with an objective-based loss that solves
@@
$\begin{aligned} &\operatorname*{minimize}\;c^{T}x\\ &\operatorname*{s.t.}\;Ax+s=b\\ &\qquad(x,s)\in\mathbb{R}^{n}\times\mathcal{K}\end{aligned}\qquad\begin{aligned} &\operatorname*{maximize}\;-b^{T}y\\ &\operatorname*{s.t.}\;-A^{T}y+r=c\\ &\qquad(r,y)\in\{0\}^{n}\times\mathcal{K}^{*}\end{aligned}$ (3.21)
where $x\in\mathbb{R}^{n}$ is the primal variable, $s\in\mathbb{R}^{m}$ is the primal slack variable, $y\in\mathbb{R}^{m}$ is the dual variable, and $r\in\mathbb{R}^{n}$ is the dual residual. The set $\mathcal{K}\in\mathbb{R}^{m}$ is a non-empty convex cone with dual cone $\mathcal{K}^{*}\in\mathbb{R}^{m}$. where $x\in\mathbb{R}^{n}$ is the primal variable, $s\in\mathbb{R}^{m}$ is the primal slack variable, $y\in\mathbb{R}^{m}$ is the dual variable, and $r\in\mathbb{R}^{n}$ is the dual residual. The set $\mathcal{K}\in\mathbb{R}^{m}$ is a non-empty convex cone. SCS uses the homogeneous self-dual embedding to view eq.˜3.21 as a fixed-point computation over ${\mathcal{Z}}=\mathbb{R}^{n\times m\times 1}$ with a scalar-valued scaling factor as the last component.
NeuralSCS proposes a semi-amortized model to predict the solution to the fixed point of the self-dual embedding that solves eq.˜3.21. Their semi-amortized model $\hat{z}_{\theta}(\phi)$ takes a context $\phi$ as the input and outputs a solution to the self-dual embedding by composing the SCS iterations $f$ with the learned fixed-point acceleration modules (${\rm init}_{\theta},{\rm acc}_{\theta}$).